From a class of girls and boys, the probability that one student chosen at random will be a boy is 2/3. If 3 girls leave the class, the probability that a student chosen at random will be a boy is 4/5. How many boys and girls are ther in the class before the 3 girls leave?
Thanks !

Answer by stanbon(75887)

From a class of girls and boys, the probability that one student chosen at random will be a boy is 2/3. If 3 girls leave the class, the probability that a student chosen at random will be a boy is 4/5. How many boys and girls are ther in the class before the 3 girls leave?
--------------------------
P(b) = 2/3
b/(b+g) = 2/3
3b = 2b+2g
b = (2/3)b + (2/3)g
(1/3)b = (2/3)g
b = 2g
------------
Equation:
b/(b + g-3) = 4/5
------
(2g)/(2g + g-3) = 4/5
10g = 4(3g-3)
10g = 12g-12
2g = 12
g = 6 (original # of girls)
----
b = 2g = 12 (original # of boys)
====================================
Cheers,
Stan H.
================